# Towards a theory of natural directed paths

### v3

@@ 4. Ziemiański’s cube chain
The “only if” direction is a consequence of the definitions of $\delta_{i,A,B}$ and $\delta_{\underline{f}}$: note that only a map of the form $\delta_{i,A,B}$ removes vertices. Let us treat the “if” part. Let $n=n_{1}=n_{2}$. Since $\operatorname{Vert}(\underline{n_{2}})\subset\operatorname{Vert}(\underline{n_{1}})$ by hypothesis, a map of $\mathcal{A}$-sets from $\mathcal{A}[\underline{n_{1}}]$ to $\mathcal{A}[\underline{n_{2}}]$ is of the form $k_{1}*\dots*k_{r}$ with $\ell(\underline{n_{2}})=r$ and where
$k_{j}:\square[\underline{m_{j}}]\longrightarrow\square[|\underline{m_{j}}|]$
is a map of $\mathcal{A}$-sets for all $1\leqslant j\leqslant r$ with $\underline{n_{1}}$ being the concatenation of the $[\underline{m_{j}}]$ for $1\leqslant j\leqslant r$ and with $\underline{n_{2}}=[|\underline{m_{1}}|,\dots,|\underline{m_{r}}|]$. We are then reduced to studying the case of a map of $\mathcal{A}$-sets
@@
[…]
@@
${\mathcal{A}[\underline{n}]\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\delta_{(g_{1},\dots,g_{p})}}$${\mathcal{A}[\underline{n}]\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathcal{L}_{\mathcal{A}}(h_{1})*\dots*\mathcal{L}_{\mathcal{A}}(h_{p})}$${\mathcal{A}[n]}$
@@
4.6 Corollary.
For any precubical $K$, Ziemiański’s definition of $\operatorname{Ch}(K)$ given in [19, Definition 1.1] or [20, Section 7] and the above definition coincide.
@@
${\mathcal{A}[m]\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{g_{1}}$${\mathcal{A}[m]\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathcal{L}_{\mathcal{A}}(k_{1})}$$\scriptstyle{\mathcal{L}_{\mathcal{A}}(h_{1})}$${\mathcal{L}_{\mathcal{A}}(K)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{A}[m]\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{g_{3}}$${\mathcal{A}[p]\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathcal{L}_{\mathcal{A}}(h_{3})}$${\mathcal{L}_{\mathcal{A}}(K)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{A}[m]\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{g_{2}}$${\mathcal{A}[m]\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathcal{L}_{\mathcal{A}}(k_{2})}$$\scriptstyle{\mathcal{L}_{\mathcal{A}}(h_{2})}$${\mathcal{L}_{\mathcal{A}}(K)}$
@@
${\mathcal{A}[\underline{n}]\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{g}$${\mathcal{A}[\underline{n}]\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathcal{L}_{\mathcal{A}}(h)}$${\mathcal{L}_{\mathcal{A}}(K)}$
@@
Let $\underline{n}=(n_{1},\dots,n_{p})\in\operatorname{Seq}(n)$. Giving $f$ is equivalent to giving $p$ maps of $\mathcal{A}$-sets $f_{i}:\mathcal{A}[n_{i}]\to\mathcal{L}_{\mathcal{A}}(K)$ for $1\leqslant i\leqslant p$ satisfying $f_{i}(1_{n_{i}})=f_{i+1}(0_{n_{i+1}})$ for $11\leqslant i\leqslant p-$ 1. Write
@@
In the same way, given $h$ is equivalent to giving $p$ maps of precubical sets $h_{i}:\square[n_{i}]\to K$ for $1\leqslant i\leqslant p$ satisfying $h_{i}(1_{n_{i}})=h_{i+1}(0_{n_{i+1}})$ for $11\leqslant i\leqslant p-$ 1. Write
@@
${\mathcal{A}[\underline{m}]\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\underline{c}}$$\scriptstyle{g_{0}}$${\mathcal{L}_{\mathcal{A}}(\square[\underline{m}])\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathcal{L}_{\mathcal{A}}(h_{1})}$$\scriptstyle{\mathcal{L}_{\mathcal{A}}(h_{0})}$${\mathcal{A}[\underline{m}]\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{g}$$\scriptstyle{\underline{c}}$${\mathcal{L}_{\mathcal{A}}(\square[\underline{n}])\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathcal{L}_{\mathcal{A}}(h)}$${\mathcal{L}_{\mathcal{A}}(K)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{L}_{\mathcal{A}}(K)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{L}_{\mathcal{A}}(K)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{L}_{\mathcal{A}}(K)}$
Figure 1. Construction of $h_{1}$
4.10 Theorem.
@@
By Quillen’s Theorem A [14, Theorem 19.6.14], it suffices to prove that the comma category $(\underline{c}{\downarrow}\mathcal{L}_{\mathcal{A}}^{K})$ is contractible for all cube chains $\underline{c}$ of $\mathcal{L}_{\mathcal{A}}(K)$. By [14, Proposition 14.3.14], it suffices to prove that the comma category $(\underline{c}{\downarrow}\mathcal{L}_{\mathcal{A}}^{K})$ has an initial object for all cube chains $\underline{c}$ of $\mathcal{L}_{\mathcal{A}}(K)$. An object of the comma category $(\underline{c}{\downarrow}\mathcal{L}_{\mathcal{A}}^{K})$ is a commutative square of $\mathcal{A}$-sets of the form
${\mathcal{A}[\underline{m}]\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\underline{c}}$$\scriptstyle{g}$${\mathcal{L}_{\mathcal{A}}(\square[\underline{n}])\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathcal{L}_{\mathcal{A}}(h)}$${\mathcal{L}_{\mathcal{A}}(K)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{L}_{\mathcal{A}}(K)}$
@@
${\mathcal{A}[\underline{m}]\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\underline{c}}$$\scriptstyle{g_{0}}$${\mathcal{L}_{\mathcal{A}}(\square[\underline{m}])\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathcal{L}_{\mathcal{A}}(h_{0})}$${\mathcal{L}_{\mathcal{A}}(K)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{L}_{\mathcal{A}}(K).}$
@@
${\mathcal{A}[\underline{m}]\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{g_{0}}$$\scriptstyle{g_{1}}$${\mathcal{L}_{\mathcal{A}}(\square[\underline{m}])\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathcal{L}_{\mathcal{A}}(h_{1})}$$\scriptstyle{\mathcal{L}_{\mathcal{A}}(h_{0})}$${\mathcal{L}_{\mathcal{A}}(K)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{A}[\underline{m}]\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{g}$${\mathcal{L}_{\mathcal{A}}(\square[\underline{n}])\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathcal{L}_{\mathcal{A}}(h)}$${\mathcal{L}_{\mathcal{A}}(K)}$
@@
${\mathcal{A}[\underline{m}]\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{g_{0}}$${\mathcal{L}_{\mathcal{A}}(\square[\underline{m}])\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathcal{L}_{\mathcal{A}}(h_{1})}$$\scriptstyle{\mathcal{L}_{\mathcal{A}}(h_{0})}$${\mathcal{L}_{\mathcal{A}}(K)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{A}[\underline{m}]\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{g}$${\mathcal{L}_{\mathcal{A}}(\square[\underline{n}])\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathcal{L}_{\mathcal{A}}(h)}$${\mathcal{L}_{\mathcal{A}}(K)}$
@@
${\square[\underline{m}]\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{h_{0}}$$\scriptstyle{h_{1}}$${\square[\underline{n}]\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{h}$${K\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${K}$
@@
${\mathcal{A}[\underline{m}]\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\underline{c}}$$\scriptstyle{g_{0}}$${\mathcal{L}_{\mathcal{A}}(\square[\underline{m}])\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathcal{L}_{\mathcal{A}}(h_{0})}$${\mathcal{L}_{\mathcal{A}}(K)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{L}_{\mathcal{A}}(K)}$
@@
4.11 Corollary.
@@
There are the homeomorphisms (natural with respect to $[n]\in\square$)
@@
the two homeomorphisms on the left by definition of the geometric realization, the right-hand homeomorphism by Proposition 2.12. Since all involved functors are colimit-preserving, we obtain for all precubical sets $K$ the natural homeomorphism
@@
Hence the space of tame natural $d$-paths of $K$ is equal to the one of $\mathcal{L}_{\mathcal{A}}(K)$. By [20, Theorem 7.5], we deduce that the space of tame natural $d$-paths of $\mathcal{L}_{\mathcal{A}}(K)$ is homotopy equivalent to $|\operatorname{Ch}(K)|$, and therefore homotopy equivalent to $|\operatorname{Ch}(\mathcal{L}_{\mathcal{A}}(K))|$ by Theorem 4.10. ∎
5. Application
We refer to [5, 6] for further details. [6, Theorem 4.1.8] states that, for two labelled cubes $\square[a_{1},\dots,a_{m}]$ and $\square[a_{m+1},\dots,a_{m+n}]$ with $m\geqslant 0$ and $n\geqslant 0$, there is an isomorphism of $\sigma$-labelled symmetric transverse sets
^$\mathcal{L}_{\widehat{\square}_{S}}\big{(}\operatorname{\overrightarrow{\operatorname{cosk}}}^{\Sigma}(\square[a_{1},\dots,a_{m}]_{\leqslant 1}\times_{\Sigma}\square[a_{m+1},\dots,a_{m+n}]_{\leqslant 1})\big{)}\\ \cong\operatorname{cosk}_{1}^{\widehat{\square}_{S},\overline{\Sigma}}\big{(}\widehat{\square}_{S}[a_{1},\dots,a_{m}]_{\leqslant 1}\overline{\times}_{\Sigma}\widehat{\square}_{S}[a_{m+1},\dots,a_{m+n}]_{\leqslant 1}\big{)}$

### v4

@@ 4 Ziemiański’s cube chain
The “only if” direction is a consequence of the definitions of $\delta_{i,A,B}$ and $\delta_{\underline{f}}$: note that only a map of the form $\delta_{i,A,B}$ removes vertices. Let us treat the “if” part. Let $n=n_{1}=n_{2}$. Write $\underline{n_{2}}=(m_{1},\dots,m_{r})$ with $\ell(\underline{n_{2}})=r$. Since $\operatorname{Vert}(\underline{n_{2}})\subset\operatorname{Vert}(\underline{n_{1}})$ by hypothesis, there exist $\underline{m_{j}}\in\operatorname{Seq}(m_{j})$ for $1\leqslant j\leqslant r$ such that $\underline{n_{1}}=(\underline{m_{1}},\dots,\underline{m_{r}})$. The smallest (biggest resp.) element of $\operatorname{Vert}(\underline{n_{2}})\subset\operatorname{Vert}(\underline{n_{1}})$ is 0 ($n$ resp.) and they can be identified with the initial and final states respectively of $\mathcal{A}[\underline{n_{1}}]$ and $\mathcal{A}[\underline{n_{2}}]$. Consider a map of $\mathcal{A}$-sets
$k:\mathcal{A}[\underline{n_{1}}]\longrightarrow\mathcal{A}[\underline{n_{2}}].$
Consider two vertices $x,y$ of $\mathcal{A}[\underline{n_{1}}]$ such that $\overrightarrow{d}\!_{1}(x,y)=1$. Then, by Proposition 4.3, the two vertices $x$ and $y$ are the extremities of a unique $1$-cube of $\mathcal{A}[\underline{n_{1}}]$ going from $x$ to $y$ which is taken by $k$ to a $1$-cube of $\mathcal{A}[\underline{n_{2}}]$. This implies that $\overrightarrow{d}\!_{1}(k(x),k(y))=1$ by Proposition 4.3 again. Consider a vertex $x$ of $\mathcal{A}[\underline{n_{1}}]$ and a sequence $(0=x_{0},\dots,x_{p}=x)$ of vertices of $\mathcal{A}[\underline{n_{1}}]$ such that $\overrightarrow{d}\!_{1}(x_{i},x_{i+1})=1$ for all $i\in\{0,\dots,p-1\}$, which implies that $\overrightarrow{d}\!_{1}(0,x)=p$ by Proposition 4.3. Then $\overrightarrow{d}\!_{1}(k(x_{i}),k(x_{i+1}))=1$ for all $i\in\{0,\dots,p-1\}$, which implies $\overrightarrow{d}\!_{1}(k(0),k(x))=p$ as well. From $\overrightarrow{d}\!_{1}(0,n)=n$ by Proposition 4.3, we then deduce that $\overrightarrow{d}\!_{1}(k(0),k(n))=n$. The only possibility is that $0=k(0)$ and that $n=k(n)$ by Proposition 4.3. We deduce
$\forall x\in\mathcal{A}[\underline{n_{1}}]_{0},\overrightarrow{d}\!_{1}(0,x)=\overrightarrow{d}\!_{1}(0,k(x)).$
Note that, in general, there is only the inequality $\overrightarrow{d}\!_{1}(x,y)\geqslant\overrightarrow{d}\!_{1}(k(x),k(y))$ for all vertices $x$ and $y$ of $\mathcal{A}[\underline{n_{1}}]$. For example, $x$ and $y$ can be incomparable vertices with $k(x)=k(y)$, which implies $\overrightarrow{d}\!_{1}(x,y)=\infty$ and $\overrightarrow{d}\!_{1}(k(x),k(y))=0$. The point is that 0 is less than all vertices of $\mathcal{A}[\underline{n_{1}}]$ for the poset structure on vertices. This implies that for all $p\in\operatorname{Vert}(\underline{n_{2}})\subset\operatorname{Vert}(\underline{n_{1}})$, $\overrightarrow{d}\!_{1}(0,p)=\overrightarrow{d}\!_{1}(0,k(p))=p$. We deduce
$\forall p\in\operatorname{Vert}(\underline{n_{2}})\subset\operatorname{Vert}(\underline{n_{1}}),k(p)=p$
because there is only one vertex $x$ in the $\underline{n_{1}}$-cube and in the $\underline{n_{2}}$-cube such that $\overrightarrow{d}\!_{1}(0,x)=p$. Thus the map
$k:\mathcal{A}[\underline{n_{1}}]\longrightarrow\mathcal{A}[\underline{n_{2}}]$
is of the form $k_{1}*\dots*k_{r}$ where
$k_{j}:\mathcal{A}[\underline{m_{j}}]\longrightarrow\mathcal{A}[|\underline{m_{j}}|]$
is a map of $\mathcal{A}$-sets for all $1\leqslant j\leqslant r$. We are then reduced to studying the case of a map of $\mathcal{A}$-sets
@@
$\mathcal{A}[n_{i}]$$\mathcal{A}[\underline{n}]$$\mathcal{A}[n]$$\mathcal{A}[n_{i}]$$\mathcal{A}[n_{i}]$$\mathcal{A}[n]$$\scriptstyle{f_{i}}\scriptstyle{\subset}$$\scriptstyle{f}$$\scriptstyle{g_{i}}$$\scriptstyle{\mathcal{L}_{\mathcal{A}}(h_{i})}$
@@
$\mathcal{A}[\underline{n}]$$\mathcal{A}[\underline{n}]$$\mathcal{A}[n]$$\scriptstyle{\delta_{(g_{1},\dots,g_{p})}}$$\scriptstyle{\mathcal{L}_{\mathcal{A}}(h_{1})*\dots*\mathcal{L}_{\mathcal{A}}(h_{p})}$
@@
Corollary 4.7.
For any precubical $K$, Ziemiański’s definition of $\operatorname{Ch}(K)$ given in (MR3722069, , Definition 1.1) or (MR4070250, , Section 7) and the above definition coincide.
@@
$\mathcal{A}[m]$$\mathcal{A}[m]$$\mathcal{L}_{\mathcal{A}}(K)$$\mathcal{A}[m]$$\mathcal{A}[p]$$\mathcal{L}_{\mathcal{A}}(K)$$\mathcal{A}[m]$$\mathcal{A}[m]$$\mathcal{L}_{\mathcal{A}}(K)$$\scriptstyle{g_{1}}$$\scriptstyle{\mathcal{L}_{\mathcal{A}}(k_{1})}$$\scriptstyle{\mathcal{L}_{\mathcal{A}}(h_{1})}$$\scriptstyle{g_{3}}$$\scriptstyle{\mathcal{L}_{\mathcal{A}}(h_{3})}$$\scriptstyle{g_{2}}$$\scriptstyle{\mathcal{L}_{\mathcal{A}}(k_{2})}$$\scriptstyle{\mathcal{L}_{\mathcal{A}}(h_{2})}$
@@
$\mathcal{A}[\underline{n}]$$\mathcal{A}[\underline{n}]$$\mathcal{L}_{\mathcal{A}}(K)$$\scriptstyle{g}$$\scriptstyle{\mathcal{L}_{\mathcal{A}}(h)}$
@@
Let $\underline{n}=(n_{1},\dots,n_{p})\in\operatorname{Seq}(n)$. Giving $f$ is equivalent to giving $p$ maps of $\mathcal{A}$-sets $f_{i}:\mathcal{A}[n_{i}]\to\mathcal{L}_{\mathcal{A}}(K)$ for $1\leqslant i\leqslant p$ satisfying $f_{i}(1_{n_{i}})=f_{i+1}(0_{n_{i+1}})$ for $1\leqslant i\leqslant p-1$. Write
@@
In the same way, given $h$ is equivalent to giving $p$ maps of precubical sets $h_{i}:\square[n_{i}]\to K$ for $1\leqslant i\leqslant p$ satisfying $h_{i}(1_{n_{i}})=h_{i+1}(0_{n_{i+1}})$ for $1\leqslant i\leqslant p-1$. Write
@@
$\mathcal{A}[\underline{m}]$$\mathcal{L}_{\mathcal{A}}(\square[\underline{m}])$$\mathcal{A}[\underline{m}]$$\mathcal{L}_{\mathcal{A}}(\square[\underline{n}])$$\mathcal{L}_{\mathcal{A}}(K)$$\mathcal{L}_{\mathcal{A}}(K)$$\mathcal{L}_{\mathcal{A}}(K)$$\mathcal{L}_{\mathcal{A}}(K)$$\scriptstyle{\underline{c}}$$\scriptstyle{g_{0}}$$\scriptstyle{\mathcal{L}_{\mathcal{A}}(h_{1})}$$\scriptstyle{\mathcal{L}_{\mathcal{A}}(h_{0})}$$\scriptstyle{g}$$\scriptstyle{\mathcal{L}_{\mathcal{A}}(h)}$$\scriptstyle{\underline{c}}$
Figure 1: Construction of $h_{1}$
Theorem 4.11.
@@
By Quillen’s Theorem A (ref_model2, , Theorem 19.6.14), it suffices to prove that the comma category $(\underline{c}{\downarrow}\mathcal{L}_{\mathcal{A}}^{K})$ is contractible for all cube chains $\underline{c}$ of $\mathcal{L}_{\mathcal{A}}(K)$. By (ref_model2, , Proposition 14.3.14), it suffices to prove that the comma category $(\underline{c}{\downarrow}\mathcal{L}_{\mathcal{A}}^{K})$ has an initial object for all cube chains $\underline{c}$ of $\mathcal{L}_{\mathcal{A}}(K)$. An object of the comma category $(\underline{c}{\downarrow}\mathcal{L}_{\mathcal{A}}^{K})$ is a commutative square of $\mathcal{A}$-sets of the form
$\mathcal{A}[\underline{m}]$$\mathcal{L}_{\mathcal{A}}(\square[\underline{n}])$$\mathcal{L}_{\mathcal{A}}(K)$$\mathcal{L}_{\mathcal{A}}(K)$$\scriptstyle{\underline{c}}$$\scriptstyle{g}$$\scriptstyle{\mathcal{L}_{\mathcal{A}}(h)}$
@@
$\mathcal{A}[\underline{m}]$$\mathcal{L}_{\mathcal{A}}(\square[\underline{m}])$$\mathcal{L}_{\mathcal{A}}(K)$$\mathcal{L}_{\mathcal{A}}(K).$$\scriptstyle{\underline{c}}$$\scriptstyle{g_{0}}$$\scriptstyle{\mathcal{L}_{\mathcal{A}}(h_{0})}$
@@
$\mathcal{A}[\underline{m}]$$\mathcal{L}_{\mathcal{A}}(\square[\underline{m}])$$\mathcal{L}_{\mathcal{A}}(K)$$\mathcal{A}[\underline{m}]$$\mathcal{L}_{\mathcal{A}}(\square[\underline{n}])$$\mathcal{L}_{\mathcal{A}}(K)$$\scriptstyle{g_{0}}$$\scriptstyle{g_{1}}$$\scriptstyle{\mathcal{L}_{\mathcal{A}}(h_{1})}$$\scriptstyle{\mathcal{L}_{\mathcal{A}}(h_{0})}$$\scriptstyle{g}$$\scriptstyle{\mathcal{L}_{\mathcal{A}}(h)}$
@@
$\mathcal{A}[\underline{m}]$$\mathcal{L}_{\mathcal{A}}(\square[\underline{m}])$$\mathcal{L}_{\mathcal{A}}(K)$$\mathcal{A}[\underline{m}]$$\mathcal{L}_{\mathcal{A}}(\square[\underline{n}])$$\mathcal{L}_{\mathcal{A}}(K)$$\scriptstyle{g_{0}}$$\scriptstyle{\mathcal{L}_{\mathcal{A}}(h_{1})}$$\scriptstyle{\mathcal{L}_{\mathcal{A}}(h_{0})}$$\scriptstyle{g}$$\scriptstyle{\mathcal{L}_{\mathcal{A}}(h)}$
@@
[…]
@@
$\mathcal{A}[\underline{m}]$$\mathcal{L}_{\mathcal{A}}(\square[\underline{m}])$$\mathcal{L}_{\mathcal{A}}(K)$$\mathcal{L}_{\mathcal{A}}(K)$$\scriptstyle{\underline{c}}$$\scriptstyle{g_{0}}$$\scriptstyle{\mathcal{L}_{\mathcal{A}}(h_{0})}$
@@
We recall (DirectedDegeneracy, , Definition 4.11) which is necessary to understand Corollary 4.12. There are the homeomorphisms (natural with respect to $n]\in\square$)
$|\square[n]|_{geom}\cong[0,1]^{n}\cong|\mathcal{A}[n]|_{geom}$
by definition of the geometric realization. A tame $d$-path of $|\mathcal{A}[n]|_{geom}=[0,1]^{n}$ is a nonconstant continuous map $\gamma:[0,\ell]\to[0,1]^{n}$ with $\ell>0$ such that $\gamma(0),\gamma(\ell)\in\{0,1\}^{n}$ and such that $\gamma$ is nondecreasing with respect to each axis of coordinates. Let $c\in K_{n}$ with $n\geqslant 1$ be an $n$-cube of an $\mathcal{A}$-set $K$. A tame $d$-path of $c$ is a composite continuous map denoted by $c;\gamma]:[0,\ell]\to|K|_{geom}$ with $\ell>0$ such that $\gamma:[0,\ell]\to[0,1]^{n}$ is a $d$-path with $c;\gamma]=|c|_{geom}\gamma$. Let $K$ be a general $\mathcal{A}$-set. A tame $d$-path of $K$ is a continuous path $0,\ell]\to|K|_{geom}$ which is a Moore composition $c_{1};\gamma_{1}]*\dots*[c_{n};\gamma_{n}]$ of $d$-paths of the cubes $c_{1},\dots,c_{n}$ of $K$. $\gamma(0)\in K_{0}$ is called the initial state of $\gamma$ and $\gamma(\ell)\in K_{0}$ is called the final state of $\gamma$.
Corollary 4.12.
@@
There are the homeomorphisms (natural with respect to $n]\in\square$)
@@
by Proposition 2.12. Since all involved functors are colimit-preserving, we obtain for all precubical sets $K$ the natural homeomorphism
@@
Hence the space of tame natural $d$-paths of $K$ is equal to the one of $\mathcal{L}_{\mathcal{A}}(K)$. By (MR4070250, , Theorem 7.5), we deduce that the space of tame natural $d$-paths of $\mathcal{L}_{\mathcal{A}}(K)$ is homotopy equivalent to $|\operatorname{Ch}(K)|$, and therefore homotopy equivalent to $|\operatorname{Ch}(\mathcal{L}_{\mathcal{A}}(K))|$ by Theorem 4.11. ∎
5 Application
We refer to ccsprecub ; symcub for further details. (symcub, , Theorem 4.1.8) states that, for two labelled cubes $\square[a_{1},\dots,a_{m}]$ and $\square[a_{m+1},\dots,a_{m+n}]$ with $m\geqslant 0$ and $n\geqslant 0$, there is an isomorphism of $\sigma$-labelled symmetric transverse sets
$\mathcal{L}_{\widehat{\square}_{S}}\big(\operatorname{\overrightarrow{\operatorname{cosk}}}^{\Sigma}(\square[a_{1},\dots,a_{m}]_{\leqslant 1}\times_{\Sigma}\square[a_{m+1},\dots,a_{m+n}]_{\leqslant 1})\big)\\ \cong\operatorname{cosk}_{1}^{\widehat{\square}_{S},\overline{\Sigma}}\big(\widehat{\square}_{S}[a_{1},\dots,a_{m}]_{\leqslant 1}\overline{\times}_{\Sigma}\widehat{\square}_{S}[a_{m+1},\dots,a_{m+n}]_{\leqslant 1}\big)$